# Reliability and Interpretability in Science and Deep Learning

### v2

@@ 4 Assumptions, simplicity and interpretability
The previous discussions started from different perspectives which all led to the same conclusion: the reliability of a model necessarily depends on the reliability of all its model assumptions, which are never self-justifying. Empirical evidence alone is never enough, neither in ML nor in TS. In Sec. 4.1 we review which classes of assumptions characterise TS models, DNN models and other ML models. While we cannot assess the reliability of any of these assumptions, in Sec. 4.2 we propose a measure of their simplicity as the only surrogate that we can access. In Sec. 4.3 we identify a close relation between the simplicity of the assumptions and the concept of interpretability (now intensively discussed within the field of responsible AI).
Any model, whether TS or ML is essentially characterised by all its assumptions and its accuracy with respect to the existing empirical data [^10] Defining a scientific or ML model by its assumptions might suggest that our discussion is restricted to a syntactic approach to scientific models. However, following (Lutz, 2017, 2014), it also applies to a semantic approach. Other views exist in the philosophical literature (Winther, 2021). But, there is currently no compelling evidence of real models that do not admit a formal description.. In the following, the expression “all assumptions” refers to all those that are necessary to reproduce the outcome of the model, including the comparison with the empirical evidence and the expression of the input data [^11] If the model predicts an outcome with some statistical distribution, as it is mostly the case, the requirement is to reproduce the statistical distribution and not the individual outcomes.. The training data themselves are not part of the model assumptions [^12] More precisely, input data are included, but they do not matter in any comparison. See 32.: they contribute to the empirical accuracy of the model just like the experiments that contribute to the development of TS models.
@@ 4.1 Assumptions
TS models rely on multiple categories of assumptions: besides the specific assumptions of the model itself, they rely also on multiple scientific disciplines that are not the main focus of the specific TS model but are essential to constrain the TS model. These disciplines are referred to as background science and they include a variety of domains from logic, mathematics, basic physics, chemistry and the modelling of any relevant experimental device. TS models typically include also assumptions to restrict their own domain of applicability [^13] Note that scope restrictions are logically equivalent to additional assumptions and are treated here as such.. In particular, TS models assume suitable values of the model parameters [^14] Alternatively, one might assume that the free parameters are determined by Maximum Likelihood principle (in a frequentist approach) or via a Bayesian argument that includes the assumption of Bayesian priors. However, this requires also the assumption of the all the empirical evidence necessary to compute the likelihood function, which is typically less concise than assuming the values of the parameters (see Sec. 4.2 and Footnote 32).. The parameters inherited from background science must be consistent with a much broader spectrum of empirical evidence from very different domains. In this sense, traditional science adopts a kind of divide et impera strategy to parameter fitting.
@@
Furthermore, TS models strive to be consistent, when applied to different data and different domains. In particular, it is typically ensured that the values of the parameters remain consistent, within expected uncertainties, across different applications. Too large deviations are interpreted as failures of the model. DNN parameters, on the contrary, are usually not regarded as something that should be consistent across applications. Although DNN training often starts from DNN models that were pre-trained on other datasets, further training is always performed for new applications and no constraint is usually imposed to keep the parameters close to the original values. That means that different assumptions are made by the DNN models for different domains of application.
@@ 4.2 Simplicity of the assumptions
The previous section shows that the key difference between TS models and DNN models is that the former use few assumptions for a wide range of phenomena. A classic scientific-philosophical tradition singles out precisely this aspect as the main non-empirical epistemic value that scientific models should try to achieve, in addition to empirical accuracy (see e.g., Galilei (1962), p. 397, Newton (1964), p. 398, Lavoisier (1862), p. 623, Poincaré (1902), Mach (1882), Weyl (1932), Barnett (1950), Kemeny (1953), Quine (1963), Lewis (1973), Walsh (1979), Derkse (1992), Swinburne (1997), Nolan (1997), Scorzato (2013) also reviewed e.g. in Baker (2004), Fitzpatrick (2014), Zellner et al. (2001)).
@@
An obvious problem with this view is that the number (and the length) of the assumptions is catastrophically language-dependent: it can always be made trivial—hence meaningless—by a suitable choice of language [^17] It is sometime noted that Kolgomorov complexity depends on the language only via a constant (Li and Vitányi, 2019). However, this is not sufficient for the purpose of model selection (Votsis, 2016).. However, it was observed in Scorzato (2013) that requiring the inclusion of a basis [^18] A basis is a set of quantities assumed to be measurable, that is sufficient to define all measurable quantities that are needed for the comparison with existing empirical evidence. of measurable concepts among the assumptions enforces a minimal achievable complexity (see details in App. A and App. B). The original argument of Scorzato (2013) is applicable to a wide class of TS models, but not immediately to DNN models. It is shown in the App. A that the mere possibility of the existence of adversarials enables the extension of the argument to DNN as well. This enables the reference to the epistemic complexity of a model in a precise and language independent way (see App. B).
@@
How does the epistemic complexity of a model impact its reliability? This question does not have a simple answer. Certainly, we cannot dismiss DNN models on the grounds of their high complexity, because we mostly do not have simpler TS models that cover the same topics and complexity-based model selection applies only to empirically equivalent models (see App. B for more details). However, a high complexity affects reliability at least in two indirect ways. The first one is interpretability. If it is challenging for any human to even understand which are all the assumptions of the model or memorise them, it becomes difficult to even formulate precisely the question: what is it whose reliability we are investigating? In this sense, a manageable epistemic complexity should be a precondition to assessing reliability. The next Sec. 4.3 looks in greater detail into the relation between the epistemic complexity introduced here and the concept of interpretability that plays a prominent role in the current discussion about responsible AI.
@@
The second indirect relation between simplicity and reliability is via (scientific) progress. It was observed already in (Popper, 1959) that simpler models offer a more direct path toward scientific progress. The present framework offers further support for this idea: simpler models provide fewer ways in which they can be changed. Hence, they point more clearly to what must be changed, or they point to a fundamental deficiency of the model. Complex models might be more flexible and survive more empirical challenges, but that is not an advantage if we are looking for a better model. As a matter of fact, TS models are the result of a long history of small and big unambiguous improvements, where different models are unified under a single one (unification) or where the parameters of some empirical model are derived from a more fundamental one (reduction) or radically new models supersede the the existing ones via revolutionary changes. This history of progresses is certainly part of their perceived reliability. On the other hand, DNN models are too complex to allow the identification of this kind of unambiguous progress: every fine tuning of paramters or architecture adjustment might bring some improvements, but it is difficult to tell what might be lost. This topic is discussed further in Sec. 5.
@@ 4.3 Interpretability
It is essential to appreciate that what matters is to understand the causes (i.e. the model assumptions) and not the entire internal mechanism that leads, step by step, from a given input to the corresponding model output. Most of our best scientific theories do not allow such “understanding”, which is often encoded in millions of operations executed by computers. Not even the best world experts would be able to reproduce the model output without the support of those computers. But understanding these details is pointless [^21] See also the discussion in Beisbart (2021) leading to similar conclusions concerning computer simulations., because all these operations are entirely determined by a few equations, which are the true causes, i.e. the model assumptions [^22] Knowledge of the model assumptions is a precondition for counterfactual explanations (Baron, 2023; Buijsman, 2022), but it is not sufficient to ensure it. Full visibility of how the output would change by any change of input may require unlimited computational power. This holds for DNN models as much as for most modern TS models.. Of course, this is true also because we know that little details, such as the specific value of the random generator used in the simulation code, do not matter. Requiring a full understanding of the detailed mechanism is an act of desperation in cases where we do not know which details actually matter, but it is not a sensible requirement when we do know that.
@@
Another interesting definition reads: “providing either an understanding of the model mechanisms and predictions, a visualisation of the model’s discrimination rules, or hints on what could perturb the model” (Arrieta et al., 2020). As mentioned above, “understanding the model mechanisms” (or visualising them) is too strong a requirement, that would unfairly penalize most modern science. However, this definition also emphasises an important aspect: control of “what can perturb the model”. As discussed in the previous section, if a set of simple rules ensures the stability of the output (not necessarily a unique exact output, but at least a unique statistical distribution which does not depend on other details), then those rules are sufficient causes/assumptions for the model’s predictions. If not, more details must be included among the causes/assumptions.
@@
Let us examine one more definition, which reads: “the degree to which a human can consistently predict the model’s result” (Kim et al., 2016). This definition tries to remove the ambiguity of the word “understanding”, which is used elsewhere, but without success: is the human allowed to utilise a tool to reproduce the model’s results? If she is, then she can just use another, identical, DNN model and every DNN model becomes perfectly interpretable. If she is not, then most modern science is not interpretable, according to this definition, because no human being can go very far without any instrument. However, this attempt contains a very interesting aspect: the concept of understanding must be clarified; we must understand enough causes (i.e. model assumptions) to reproduce the outcome, and the outcome must be “consistent”, i.e. robust to changing details. This is achieved if we know all the model assumptions, if these do not require the inclusion of too many details and if the output is computable for the relevant inputs (i.e. the input is letigimete for the model). So, also this last definition supports an identification of the concept of interpretability with a measure of the simplicity of all the assumptions.
To conclude this section, the reliability of a model crucially depends on its interpretability and the latter can be given a clear meaning that doesn’t depend on the vague notion of “understanding”, doesn’t depend on any language and doesn’t depend on individual skills.
@@ 5 Conclusions and way forward
DNNs are being used effectively in many contexts where it is not necessary to assess their reliability precisely. This is the case when DNN models generate candidate solutions that are eventually checked via independent tests (Jumper et al., 2021), or when they generate data that do not need to be sampled with a strictly defined distribution, because the conclusion has no strict statistical value or because the distribution is corrected afterwards (Albergo et al., 2019) [^23] If the independent check is performed by humans, it becomes important to compare the reliability of DNN with the one of humans. This a very important and interesting question, but beyond the scope of this paper, that focuses on comparing DNN models to TS models.
@@
DNNs often display impressive predictive power, which can be possible only because, out of their vast amount of parameters, only a few combinations can be actually relevant. However, while the existence of a submanifold of relevant parameters (and the ability of the algorithm to find it, somehow [^24] The observation that DNN seem to be able to identify the right features, sometimes, is equivalent to state that it seems to solve the reference class problem, sometimes (Buchholz, 2023). However, our current problem is to characterise sometimes.) is sufficient for their powerful predictivity, it is not sufficient for their reliability. The latter relies also on actually identifying that manifold of relevant parameters. This is necessary for interpretability, which is not just a desirable property for non-expert users: it is a precondition to assess the reliability itself. Identifying the assumptions is also necessary for scientific progress (Scorzato, 2016), because we need to know which assumptions we may want to replace.
Impressive predictive power in new domains, combined with weak foundations and difficult synthesis with background-science can be a signal of an ongoing (and incomplete) scientific revolution that may lead to a paradigm change (Kuhn, 1962), as was the case for the birth of Quantum Mechanics. What could be the way forward? Predicting the retirement of the scientific method itself (towards something like a theory-free science) appears misguided, according to the present analysis.
@@
One interesting alternative approach is to use DNN as a tool to suggest features for other models, along the lines of (Huang and LeCun, 2006; Notley and Magdon-Ismail, 2018). However, it is important to ensure that the features obtained through this process are directly measurable and not simply the output of the last DNN layer, otherwise the DNN component cannot be removed from the formal counting of the assumptions needed to produce the results. The features extracted in this way would ease the connection with background science (imposing valuable constraints) and enable the building of simpler models, possibly unrelated to DNN models.
@@
Finally, another promising avenue of research is to study those limits that can be computed exactly, as the limit of infinitely large width of fully connected DNNs (Jacot et al., 2018). Such analytical results are essential to test the behaviour of a DNN where we know exactly how it should behave. So, their first epistemological advantage is to increase the empirical accuracy of the DNNs (or rule them out). Furthermore, these limit cases could potentially also suggest how to realise simpler ML methods that are as powerful as DNN under specific circumstances. Another possibility is to use exactly solved limit cases as a starting point to set up DNNs that are small deviations from those limit cases, which might potentially enable a simpler formulation (with fewer assumptions) of the DNN itself.

### v3

@@ 4 Assumptions, simplicity and interpretability
The previous discussions started from different perspectives which all led to the same conclusion: the reliability of a model necessarily depends on the reliability of all its model assumptions, which are never self-justifying. Empirical evidence alone is never enough, neither in ML nor in TS. In Sec. 4.1 we review which classes of assumptions characterise TS models, DNN models and other ML models. While we cannot assess the reliability of these assumptions, in Sec. 4.2 we propose a measure of their simplicity as the most relevant surrogate used in the scientific practice (often implicitly). In Sec. 4.3 we identify a close relation between the simplicity of the assumptions and the concept of interpretability (now intensively discussed within the field of responsible AI).
Any model, whether TS or ML is essentially characterised by all its assumptions and its accuracy with respect to the existing empirical data [^12] Defining a scientific or ML model by its assumptions might suggest that our discussion is restricted to a syntactic approach to scientific models. However, following (Lutz, 2017, 2014), it also applies to a semantic approach. Other views exist in the philosophical literature (Winther, 2021). But, there is currently no compelling evidence of real models that do not admit a formal description.. In the following, the expression “all assumptions” refers to all those that are necessary to reproduce the outcome of the model, including the comparison with the empirical evidence and the expression of the input data [^13] If the model predicts an outcome with some statistical distribution, as it is mostly the case, the requirement is to reproduce the statistical distribution and not the individual outcomes.. Normally, the training data themselves are not part of the model assumptions [^14] Note that the input data, instead, are part of the assumptions, but they do not matter when comparing models on the same questions. See Footnote 37.: they contribute to select the model parameters (which are part of the assumptions) and they contribute to the empirical accuracy of the model just like the experiments that contribute to the development of TS models. However, including the training data among the assumptions might lead to simpler formulations of some DDN models characterised by a huge number of parameters. Hence, this possibility is also considered.
@@ 4.1 Assumptions
TS models rely on multiple categories of assumptions: besides the specific assumptions of the model itself, they rely also on multiple scientific disciplines that are not the main focus of the specific TS model but are essential to constrain the TS model. These disciplines are referred to as background science and they include a variety of domains from logic, mathematics, basic physics, chemistry and the modelling of any relevant experimental device. TS models typically include also assumptions to restrict their own domain of applicability [^15] Note that scope restrictions can be expressed as part of the assumptions (namely relativizations). Therefore, we will not discuss scope restrictions separately in this paper.. In particular, TS models assume suitable values of the model parameters [^16] Alternatively, one might assume that the free parameters are determined by Maximum Likelihood principle (in a frequentist approach) or via a Bayesian argument that includes the assumption of Bayesian priors. However, this requires also the assumption of the all the empirical evidence necessary to compute the likelihood function, which is typically less concise than assuming the values of the parameters (see Sec. 4.2 and Footnote 37).. The parameters inherited from background science must be consistent with a much broader spectrum of empirical evidence from very different domains. In this sense, traditional science adopts a kind of divide et impera strategy to parameter fitting.
@@
Furthermore, TS models strive to be consistent, when applied to different data and different domains. In particular, it is typically ensured that the values of the parameters with identical meaning remain consistent, within expected uncertainties, across different applications. Too large deviations are interpreted as failures of the model. DNN parameters, on the contrary, are usually not regarded as something that should be consistent across applications. Although DNN training often starts from DNN models that were pre-trained on other datasets, further training is always performed for new applications and no constraint is usually imposed to keep the parameters close to the original values. That means that different assumptions are made by the DNN models for different domains of application.
@@ 4.2 Simplicity of the assumptions
The previous section shows that a key difference between TS models and DNN models is that the former use relatively few assumptions and few parameters for a wide range of phenomena. A classic scientific-philosophical tradition singles out precisely this aspect as the main non-empirical epistemic value that scientific models should try to achieve, in addition to empirical accuracy (see e.g., Galilei (1962), p. 397, Newton (1964), p. 398, Lavoisier (1862), p. 623, Poincaré (1902), Mach (1882), Weyl (1932), Barnett (1950), Kemeny (1953), Quine (1963), Lewis (1973), Chaitin (1975), Zenil (2020b), Walsh (1979), Derkse (1992), Swinburne (1997), Nolan (1997), Scorzato (2013), also reviewed e.g. in Baker (2004), Fitzpatrick (2014), Zellner et al. (2001)). The first goal of this section is to clarify the meaning of “few assumptions”.
@@
An apparent problem with this view is that the number (and the length) of the assumptions is catastrophically language-dependent: it can always be made trivial—hence meaningless—by a suitable choice of language [^19] It is sometime noted that Kolmogorov complexity depends on the language only via a constant (Li and Vitányi, 2019). However, this is not sufficient for the purpose of model selection (Votsis, 2016).. However, it was observed in Scorzato (2013) that requiring the inclusion of a basis [^20] A basis is a set of quantities assumed to be measurable, that is sufficient to define all measurable quantities that are needed for the comparison with existing empirical evidence. of measurable concepts among the assumptions enforces a minimal achievable complexity (see details in App. A and App. B). The original argument of Scorzato (2013) is applicable to a wide class of TS models, but not immediately to DNN models. It is shown in the App. A that the mere possibility of the existence of adversarials enables the extension of the argument to DNN as well. This enables the reference to the epistemic complexity of a model in a way that is precise and language independent [^21] Epistemic complexity depends, of course, on the model and hence on its empirical content, which is not language dependent, but needs a language to be described (see App. B)..
@@
How does the epistemic complexity of a model impact its reliability? This question does not have a simple answer. Certainly, we cannot dismiss DNN models on the grounds of their high complexity, because we mostly do not have simpler TS models that cover the same topics and complexity-based model selection applies only to empirically equivalent models (see App. B for more details). However, a high complexity affects reliability at least in three indirect ways. The first one is interpretability. If it is challenging for any human to even understand which are all the assumptions of the model or review them, it becomes difficult to even formulate precisely the question: what is it whose reliability we are investigating? In this sense, a manageable epistemic complexity should be a precondition for a plausible assessment of reliability. The next Sec. 4.3 looks in greater detail into the relation between the epistemic complexity introduced here and the concept of interpretability that plays a prominent role in the current discussion about responsible AI.
The second relation between simplicity and reliability comes from the fact that simplicity enables the definition of the state-of-the-art models (see App. B), namely those models that represent all the possible compromises between simplicity and the many dimensions of accuracy. The state-of-the-art, in turns, represents a non-arbitrary class of models that offers a reference for a Bayesian framework for error estimation [^22] In the scientific practice, Bayesian estimates may not include the entire state-of-the-art, but the existence of the state-of-the-art justifies other choices as approximations of a well defined class of models.. The high epistemic complexity of DNN models makes it very difficult to identify the state-of-the-art. The computational cost of the Bayesian approach applied to DNNs has been widely acknowledged (Gawlikowski et al., 2021; Abdar et al., 2021; Loftus et al., 2022), but it is important to emphasise that this makes it also difficult to define the Bayesian errors in the first place.
@@
The third indirect relation between simplicity and reliability is via (scientific) progress. It was observed already in (Popper, 1959) that simpler models offer a more direct path toward scientific progress. The present framework offers further support for this idea: simpler models offer fewer opportunities for small changes. Hence, they point more clearly to what must be changed, or they point to a fundamental deficiency of the model. Complex models might be more flexible and survive more empirical challenges, but that is not an advantage if we are looking for a better model. As a matter of fact, TS models are the result of a long history of small and big unambiguous improvements, where different models are unified under a single one (unification) or where the parameters of some empirical model are derived from a more fundamental one (reduction) or radically new models supersede the the existing ones via revolutionary changes. This history of progresses is certainly part of their perceived reliability. On the other hand, DNN models are too complex to allow the identification of this kind of unambiguous progress: every fine tuning of paramters or architecture adjustment might bring some improvements, but it is difficult to tell what might be lost. This topic is discussed further in Sec. 5.
@@ 4.3 Interpretability
It is essential to appreciate that what matters is to understand the causes (i.e. the model assumptions) and not the entire internal mechanism that leads, step by step, from a given input to the corresponding model output. Most of our best scientific theories do not allow such “understanding”, which is often encoded in millions of operations executed by computers. Not even the best world experts would be able to reproduce most of the model output without the support of those computers. But understanding these details is unnecessary [^25] See also the discussion in Beisbart (2021) leading to similar conclusions concerning computer simulations., because all these operations are entirely determined by a few equations, which are the true causes, i.e. the model assumptions [^26] Knowledge of the model assumptions is a precondition for counterfactual explanations (Baron, 2023; Buijsman, 2022), but it is not sufficient to ensure it. Full visibility of how the output would change by any change of input may require unlimited computational power. This holds for DNN models as much as for most modern TS models.. Of course, this is true also because we know that little details, such as the specific value of the random generator used in the simulation code, do not matter. Requiring a full understanding of the detailed mechanism is an act of desperation in cases where we do not know which details actually matter, but it is not a sensible requirement when we do know that.
The view described above is consistent with the one defended by de Regt and Dieks (2005), who criticize the causal-mechanical conception of explanatory understanding for similar reasons to those given here. According to de Regt and Dieks (2005), experts are often able to connect some changes in model assumptions to some changes in model outcomes, without performing caluculations because they extrapolate from other results that they know. But, for most modern scientific models, even the best experts cannot predict most of the outcomes without the help of external computational tool.
@@
Another often-cited definition reads: “providing either an understanding of the model mechanisms and predictions, a visualisation of the model’s discrimination rules, or hints on what could perturb the model” (Arrieta et al., 2020). As mentioned above, “understanding the model mechanisms” (or visualising them) is too strong a requirement, that would unfairly penalize most modern science. However, this definition also emphasises an important aspect: control of “what can perturb the model”. As discussed in the previous section, if a set of simple rules ensures the stability of the output (not necessarily a unique exact output, but at least a unique statistical distribution which does not depend on other details), then those rules are sufficient causes/assumptions for the model’s predictions. If not, more details must be included among the causes/assumptions.
@@
Let us examine one more definition, which reads: “a method is interpretable if a user can correctly and efficiently predict the method’s results” (Kim et al., 2016). This definition tries to remove the ambiguity of the word “understanding”, which is used elsewhere, but without success: is the user allowed to utilise a tool to reproduce the model’s results? If she is, then she can just use another, identical, DNN model and every DNN model becomes perfectly interpretable. If she is not, then most modern science is not interpretable, according to this definition, because no human being can go very far without any instrument. This definition is not ambiguous in the context of image recognition, where it has been introduced (Kim et al., 2016), because it assumes that human vision is the tool available to the user. But, it cannot be extended to a general context, as suggested in (Molnar, 2020). However, this attempt contains a very interesting aspect: the concept of understanding must be clarified; we must understand enough causes (i.e. model assumptions) to reproduce the outcome, and the outcome must be “consistent”, i.e. robust to changing details. This is achieved if we know all the model assumptions, if these do not require the inclusion of too many details and if the output is computable for the relevant inputs (i.e. the input is legitimate for the model). So, also this last definition supports an identification of the concept of interpretability with a measure of the simplicity of all the assumptions.
In conclusion, we can attribute a precise meaning to the concept of “interpretability”. In this way, “interpretability” does not depend on the vague notion of “understanding”, does not depend arbitrarily on the choice of language, and does not depend on the individual skills of the persons interacting with the model.
Moreover, interpretability is not only a desirable property for non-expert, it is a precondition for a plausible assessment of the reliability. More precisely, lower interpretability makes the assessment of reliability less plausible. This claim can be justified as follows. Assessing the reliability of any model is done in two steps. The first step consists in identifying a class ($C$) of models that should be used to assess the reliability of the given target model (see Sec. 3.1). The second step consists in computing the confidence ranges (e.g. via the Bayesian priors and the Bayesian formula). The first step is the focus of this paper. In TS, scientists tend to agree, de facto, on which class $C$ should be used (which we can call the state-of-the-art models, for which Scorzato (2016) proposes a characterization. See Def. 3 in App. B). Whether we adopt Def. 3 or not, identifying the class $C$ requires examining non-empirical epistemic properties of the corresponding models. If we do adopt Def. 3, the state-of-the-art becomes broader when more complex models are involved, because it is increasingly more difficult to check if any of the many variations of a complex model should or should not be included in the class $C$. In fact, complex models leave larger opportunities for models that are not more complex. This leads to larger confidence ranges and larger uncertainties on the confidence ranges themselves. Even if we did not adopt Def. 3, we should still evaluate some (unclear) non-empirical properties of the models, which must include also reviewing their assumptions. The lack of precise requirements cannot simplify the assessment of more complex assumptions. This forces, again, larger uncertainties on the confidence ranges. In both cases, more complex (i.e. less interpretable) assumptions lead to less plausible assessments of reliability [^27] Note that it is common practice to penalize complex assumptions by reweighting the priors within a Bayesian framework (Schwarz, 1978). But it is also often claimed that the effect of the priors vanishes for large statistics. Here we discuss uncertainties that are not eliminated statistically..
@@ 5 Conclusions and way forward
DNNs are being used effectively in many contexts where it is not necessary to assess their reliability precisely. This is the case when DNN models generate candidate solutions that are eventually checked via independent tests (Jumper et al., 2021), or when they generate data that do not need to be sampled with a strictly defined distribution, because the conclusion has no strict statistical value or because the distribution is corrected afterwards (Albergo et al., 2019) [^28] If the independent check is performed by humans, it becomes important to compare the reliability of DNN with the one of humans. This a very important and interesting question, but beyond the scope of this paper, that focuses on comparing DNN models to TS models..
@@
DNNs often display impressive predictive power, which can only be possible because, somehow, the training process finds a DNN configuration that is sensitive to the relevant features of the data and is insensitive to the irrelevant ones [^29] The observation that DNN seem to be able to identify the right features, sometimes, is equivalent to state that it seems to solve the reference class problem, sometimes (Buchholz, 2023). However, our current problem is to characterise sometimes.. However, while the existence of such configurations (and the ability of the algorithm to find them) is sufficient for their powerful predictivity, it is not sufficient for their reliability. To define what we mean by reliability, we must clarify the class of possible alternative models (see Sec. 3.1). TS models can typically refer to the class of models that belong to the state-of-the-art or some approximations of it. Identifying the state-of-the-art is much more difficult for DNN models because of their high epistemic complexity, which makes it difficult not only to compute the Bayesian errors, but also to define them.
This paper has proposed to identify the interpretability of a model with its inverse epistemic complexity. In this sense, interpretability is not only a desirable property for non-expert users: it is a precondition for a plausible assessment of the reliability itself. Identifying the assumptions is also necessary for scientific progress (Scorzato, 2016), because we need to know which assumptions we may want to replace.
Impressive predictive power in new domains, combined with weak foundations and difficult synthesis with background-science can be a signal of an ongoing (and incomplete) scientific revolution, as was the case for the birth of Quantum Mechanics. What could be the way forward? Predicting the retirement of the scientific method itself (towards something like a theory-free science) appears misguided, according to the present analysis.
@@
One interesting alternative approach is to use DNN as a tool to suggest features for other models, along the lines of (Huang and LeCun, 2006; Notley and Magdon-Ismail, 2018). However, it is important to ensure that the features obtained through this process are directly measurable and not simply the output of the last DNN layer, otherwise the DNN component cannot be removed from the formal counting of the assumptions needed to produce the results. The features extracted in this way would ease the connection with background science (imposing valuable constraints) and enable the building of simpler models, possibly unrelated to DNN models. See (Guth et al., 2023) and references therein for significant progresses in this direction.
@@
Another promising avenue of research is to study those limits that can be computed exactly, as the limit of infinitely large width of fully connected DNNs (Jacot et al., 2018). Such analytical results are essential to test the behaviour of a DNN where we know exactly how it should behave. So, their first epistemological advantage is to increase the empirical accuracy of the DNNs (or rule them out). Furthermore, these limit cases could potentially also suggest how to realise simpler ML methods that are as powerful as DNN under specific circumstances. Another possibility is to use exactly solved limit cases as a starting point to set up DNNs that are small deviations from those limit cases, which might potentially enable a simpler formulation (with fewer assumptions) of the DNN itself.